# Towards Reliable Latent Knowledge Estimation in LLMs: In-Context Learning vs. Prompting Based Factual Knowledge Extraction

### v1

@@ Abstract
We propose an approach for estimating the latent knowledge embedded inside large language models (LLMs). We leverage the in-context learning (ICL) abilities of LLMs to estimate the extent to which an LLM knows the facts stored in a knowledge base. Our knowledge estimator avoids reliability concerns with previous prompting-based methods, is both conceptually simpler and easier to apply, and we demonstrate that it can surface more of the latent knowledge embedded in LLMs. We also investigate how different design choices affect the performance of ICL-based knowledge estimation. Using the proposed estimator, we perform a large-scale evaluation of the factual knowledge of a variety of open source LLMs, like OPT, Pythia, Llama(2), Mistral, Gemma, etc. over a large set of relations and facts from the Wikidata knowledge base. We observe differences in the factual knowledge between different model families and models of different sizes, that some relations are consistently better known than others but that models differ in the precise facts they know, and differences in the knowledge of base models and their finetuned counterparts.
Towards Reliable Latent Knowledge Estimation in LLMs:
In-Context Learning vs. Prompting Based Factual Knowledge Extraction
Qinyuan Wu1, Mohammad Aflah Khan1, Soumi Das1, Vedant Nanda1,2, Bishwamittra Ghosh1, Camila Kolling1, Till Speicher1, Laurent Bindschaedler 1, Krishna P. Gummadi1, Evimaria Terzi3, 1Max Planck Institute for Software Systems(MPI-SWS), 2University of Maryland, College Park, 3Boston University, Boston University, Correspondence: qwu@mpi-sws.org
1 Introduction
@@
Conversational chatbots (e.g., OpenAI’s ChatGPT) built around large language models (e.g., OpenAI’s GPT) are increasingly being used for a variety of information retrieval tasks such as searching for information or seeking recommendations related to real world entities like people or places (Wu et al., 2023; Zhu et al., 2023). A worrisome concern in such scenarios is the factual correctness of information generated by the LLMs (Peng et al., 2023; Hu et al., 2023a; Snyder et al., 2023; Yao et al., 2023; Ji et al., 2023; Zhang et al., 2023; Wang et al., 2023).
@@
The latent knowledge estimation problem: To avoid making false assertions about a real-world entity, an LLM first needs to have factual (true) knowledge about the entity. Given a prompt like “Einstein was born in the year”, LLMs may generate both the correct answer (“1879”) and wrong answers (e.g., “1878” or “1880”) with some probabilities. If an LLM knows the fact, one can hope that the probability with which it would generate the correct answer would be much higher than the wrong answers Jiang et al. (2021). As LLMs are typically pretrained over a Web corpus (including Wikipedia data) with millions of facts about real-world entities, they have the opportunity to learn factual knowledge about our world and latently embed the knowledge in their parameters. But, how can we estimate the extent to which LLMs have knowledge of real-world facts?
@@
Reliability of latent knowledge estimates: Prior works Jiang et al. (2020); Bouraoui et al. (2020) followed Petroni et al. (2019), and represented factual knowledge in the form of triplets , where the subject has a relation of type with the object (e.g., ). The central challenge of latent knowledge estimation is to infer given and by only using information extracted from the LLM. Typically, the inference relies on probing the LLM with prompts constructed using and and analyzing the responses. Current approaches have few well-defined rules to avoid prompt engineering and prompt hacking, raising serious concerns about the reliability of their estimates. Against this background, in this paper, we make four primary contributions:
1. A simple yet reliable latent knowledge estimator (LKE) leveraging in-context learning (ICL): We propose a latent knowledge estimator (LKE) that leverages in-context learning (ICL), called IC-LKE, in a simple yet clever way to avoid the many reliability concerns with prompting based previous knowledge estimators.
2. Exploring the nuances of using ICL for knowledge estimation: We investigate the impact of different ICL design choices on the estimation of latent knowledge, such as the number of in-context examples, when some of the examples are unknown to the model or simply incorrect, as well as the sequence in which they appear. While we focus on knowledge estimation, our findings can inform the application of ICL in other contexts.
3. A comparison of IC-LKE with previous approaches: We empirically demonstrate that IC-LKE outperforms previous knowledge estimation approaches that rely on human-generated or machine-mined prompts across a variety of different open-source models and different types of factual relations. In contrast to prompting based methods, which are relation-specific and LLM-specific, IC-LKE’s design is straightforward to apply.
@@
4. A systematic comparison of latent knowledge of open source LLMs at scale: We use IC-LKE to evaluate the knowledge of 49 open-source LLMs spanning many families such as Llama(2), Gemma, Mistral, OPT, Pythia, etc. across a wide range of sizes, both with and without instruction-finetuning over 50 different relations and 20,000 facts from Wikidata. We find that models from some families such as Llama2, Mistral and Gemma and larger models know more facts than others, that models within the same family differ in the specific facts they know, despite being trained on the same data, and that fine-tuning reduces the amount of factual knowledge that can be extracted from the models.
Related Work:
Researchers have proposed several approaches to estimate latent knowledge from LLMs, which can be categorized into two ways: (i) Model-internals based approaches leverage the LLM attention map Wang et al. (2020), activation function Burns et al. (2022), or model parameters Kazemnejad et al. (2023) to decide whether factual information can be extracted from the LLM. In our study, we rely on the probability distribution of generated tokens in an LLM – thereby our method belongs to the model-responses based approach. (ii) Model-responses based approaches – generally applicable to a wide range of LLM models – often propose different prompting techniques to nudge the LLM to validate whether a target fact is stored in it Chern et al. (2023); Sun et al. (2023); Wang et al. (2020); Petroni et al. (2019); Jiang et al. (2021); Newman et al. (2022); Jiang et al. (2020). Prompt-based methods differ subtly by the choice of prompts and evaluation criteria. Besides, the prompts are often brittle Zamfirescu-Pereira et al. (2023); Arora et al. (2023); Sclar et al. (2023) – their success depends on the hypothesis that the LLM indeed understands the prompts. In our study, we instead seek a minimal understanding of prompts by an LLM and design a knowledge estimation method based on the in-context learning. As a test bed Elsahar et al. (2018); Hu et al. (2023b); Sun et al. (2023); Petroni et al. (2019); Zhu and Li (2023); Kryściński et al. (2019), we consider facts from existing knowledge graphs for performing knowledge estimation of LLMs.
2 Designing Reliable LKEs
@@
Today, there exist many general-purpose as well as domain-specific factual knowledge bases that contain a very large number (millions to billions) of facts. The facts can be encapsulated as triplets, represented as subject (), relation (), object (). These triplets offer a general way to represent factual knowledge about real-world entities in knowledge graphs or other structured knowledge bases. The goal of latent knowledge estimation is to infer what fraction of the facts are known to a LLM. We call methods that estimate the amount of latent knowledge inside an LLM latent knowledge estimators (LKEs).
2.1 Reliability concerns with existing LKEs
@@
Existing approaches to estimating latent knowledge in LLMs use a variety of factual knowledge tests. Below, we identify several reliability concerns with current designs that motivate our new LKE design.
1. LLM-specific restrictions on test topics: Many prior works Petroni et al. (2019); Jiang et al. (2020) limit the choice of facts that can be used in tests to those where the surface form of the objects () is represented by a single token by the LLM’s tokenizer. As different LLMs use different tokenizers, this limitation prevents us from comparing the latent knowledge across different LLMs. Furthermore, only popular objects tend to be represented by a single token and so the resulting estimates are not representative of the LLM’s knowledge of facts with multi-token object representations.
@@
2. Unrestricted choice of test prompts: Many past works have attempted to use test prompts without any restrictions, including both human-generated or machine-mined prompts (Jiang et al., 2020; Zamfirescu-Pereira et al., 2023; Arora et al., 2023; Sclar et al., 2023). They typically intersperse the subject and object between additional relationship context-communicating tokens. Some analyze the performance of a variety of prompts and then pick the best-performing or use an ensemble of the best-performing prompts (Jiang et al., 2020; Newman et al., 2022; Fernando et al., 2023). However, these approaches raise two important concerns: First, the generated prompts, particularly those that are machine-mined, may include tokens that can implicitly or explicitly introduce additional (side-channel) information that makes it easier to answer the question. As a specific example, in a prior work Jiang et al. (2020), for the relation “position held", the prompt “ has the position of " performed worse than “ is elected ". But, note that the second prompt potentially introduces a side-channel: it implicitly rules out answer choices for unelected positions like Professor and favors elected positions like President. Second, selecting from an unbounded number of potential prompt choices raises concerns about the complexity of LKEs (the size of the set of all considered prompts) and the potential for over-fitting, which in turn brings the reliability of estimates into question.
@@
3. Reliance on LLMs’ meta-linguistic judgments: Prior works used prompts Chern et al. (2023); Sun et al. (2023); Wang et al. (2020); Petroni et al. (2019); Jiang et al. (2021); Newman et al. (2022); Jiang et al. (2020) for communicating the question as well as the expected format of answers. But, the scores (estimates) resulting from such prompt-based testing conflate an LLM’s latent knowledge of the facts with the LLM’s meta-linguistic judgments, i.e., the LLM’s ability to comprehend the prompt, understand the question embedded within the prompt and output the answer in some expected format (Hu and Levy, 2023). The impact on meta-linguistic judgments can be seen from the fact that multiple semantically-equivalent prompts result in different responses from an LLM and thereby, different estimates of latent knowledge (Hu and Levy, 2023).
@@
Motivated from the above, we derive the following three design principles for LKEs. A reliable LKE design should:
DP1: generate estimates for any factual topic and tokenization scheme.
@@
DP2: limit arbitrary prompt engineering to minimize over-fitting & side-channels.
DP3: minimize reliance on meta-linguistic prompts.
2.2 A new In Context learning based LKE (IC-LKE)

### v2

# Towards Reliable Latent Knowledge Estimation in LLMs: Zero-Prompt Many-Shot Based Factual Knowledge Extraction

@@ Abstract.
In this paper, we focus on the challenging task of reliably estimating factual knowledge that is embedded inside large language models (LLMs). To avoid reliability concerns with prior approaches, we propose to eliminate prompt engineering when probing LLMs for factual knowledge. Our approach, called Zero-Prompt Latent Knowledge Estimator (ZP-LKE), leverages the in-context learning ability of LLMs to communicate both the factual knowledge question as well as the expected answer format. Our knowledge estimator is both conceptually simpler (i.e., doesn’t depend on meta-linguistic judgments of LLMs) and easier to apply (i.e., is not LLM-specific), and we demonstrate that it can surface more of the latent knowledge embedded in LLMs. We also investigate how different design choices affect the performance of ZP-LKE. Using the proposed estimator, we perform a large-scale evaluation of the factual knowledge of a variety of open-source LLMs, like OPT, Pythia, Llama(2), Mistral, Gemma, etc. over a large set of relations and facts from the Wikidata knowledge base. We observe differences in the factual knowledge between different model families and models of different sizes, that some relations are consistently better known than others but that models differ in the precise facts they know, and differences in the knowledge of base models and their finetuned counterparts. 111Code available at: https://github.com/QinyuanWu0710/ZeroPrompt_LKE
1. Introduction
@@
Conversational chatbots (e.g., OpenAI’s ChatGPT) built around large language models (e.g., OpenAI’s GPT) are increasingly being used for a variety of information retrieval tasks such as searching for information or seeking recommendations related to real-world entities like people or places (Wu et al., 2023; Zhu et al., 2023). A worrisome concern in such scenarios is the factual correctness of information generated by the LLMs (Peng et al., 2023; Hu et al., 2023b; Snyder et al., 2023; Yao et al., 2023; Ji et al., 2023; Zhang et al., 2023; Wang et al., 2023; Kalo and Fichtel, [n. d.]; Cao et al., 2021; Youssef et al., 2023).
@@
The latent knowledge estimation problem: To avoid making false assertions about a real-world entity, an LLM first needs to have factual (true) knowledge about the entity. Given a prompt like “Einstein was born in the year”, LLMs may generate both the correct answer (“1879”) and wrong answers (e.g., “1878” or “1880”) with some probabilities. If an LLM knows the fact, one can hope that the probability with which it would generate the correct answer would be much higher than the wrong answers (Jiang et al., 2021). As LLMs are typically pretrained over a Web corpus (including Wikipedia data) with millions of facts about real-world entities, they have the opportunity to learn factual knowledge about our world and latently embed the knowledge in their parameters. But, how can we estimate the extent of LLMs’ knowledge of real-world facts?
@@
Reliability of latent knowledge estimates: Following (Petroni et al., 2019), many prior works (Jiang et al., 2020; Bouraoui et al., 2020; Shin et al., 2020) represent factual knowledge in the form of triplets , where the subject has a relation of type with the object (e.g., ). The central challenge of latent knowledge estimation is to infer given and by only using information extracted from the LLM. Typically, the inference relies on probing the LLM with prompt templates, , constructed to communicate the information of and and analyzing the generated responses for presence of (see Figure 1). Current approaches (Hu et al., 2023a; Kalo and Fichtel, [n. d.]; Jiang et al., 2021, 2020; Shin et al., 2020; Sun et al., 2023) allow unrestricted choice of prompt templates with few well-defined rules (see Figure 2). As a result, they are vulnerable to prompt engineering and prompt hacking, which raises serious concerns about the reliability of their estimates (Cao et al., 2021). Against this background, in this paper, we make four primary contributions:
1. A structured reliable latent knowledge estimator (LKE) based on zero prompting (ZP): Our latent knowledge estimator, called ZP-LKE, is based on the following simple, yet novel and powerful insight. Rather than engineer input prompts that best communicate to an LLM, we let the LLM infer by simply providing multiple examples of pairs that share the same relation (see Figure 2). The key distinguishing feature of ZP-LKE is its adherence to zero prompting, i.e., the input is highly structured and contains no prompt tokens outside other similarly related pairs. Thus, ZP-LKE avoids the reliability risks associated with prompt engineering such as side-channels and over-fitting. (We discuss these reliability concerns in Section 2.1.)
| Prompt based LKEs (ZS-LKE & FS-LKE) | Zero-Prompt based LKE (ZP-LKE) | ||
| Input | Output: Next tokens | Input | Output: Next tokens |
| Peter Grünberg was born in | Munich in 1939 and obtained | Albert Einstein 1879, Peter Grünberg | 2007 Mohammed Hanif |
| Albert Einstein was born in 1879, Brian Kobilka was born in 1955, Peter Grünberg was born in | 1939 and has a birthday | Albert Einstein 1879, Brian Kobilka 1955, Stefan W. Hell 1962, Ivan Pavlov 1849, Peter Grünberg | 1939, Albert Szent- |
| Albert Einstein was born in 1872, Brian Kobilka was born in 1965, Peter Grünberg was born in | 1939 and has a birthday | Albert Einstein 1872, Brian Kobilka 1927, Stefan W. Hell 1962, Ivan Pavlov 1878, Peter Grünberg | 2007, Albert Szent- |
| Author-1 was born in 1872, Author-2 was born in 1965, Peter Grünberg was born in | 1939 and has a life expect | Author-2 1879, Author-3 1955, Author-4 1962, Author-1 1849, Peter Grünberg | 2022, Rossitza |
2. ZP-LKE requires many-shots and is fundamentally different from few-shot prompting: A recent work (Kalo and Fichtel, [n. d.]) shows that few-shot prompting (FS-LKE in Figure 2) can yield improved knowledge estimates compared to zero-shot prompting (ZS-LKE in Figure 2) by providing the LLM with examples to format generated answers. In contrast, ZP-LKE also uses examples to effectively communicate the question at hand. The different modes in which examples are being used in FS-LKE and ZP-LKE is illustrated further in Table 1. Adding a few examples to the zero-shot prompt ”Peter Grünberg was born in” results in the model generating the correct answer ”1939” right-away. However, it does not appear to matter whether the examples provided are reflecting correct information or information about subjects known to the LLM (consistent with our hypothesis that the examples are being used to infer answer format). In contrast, Table 1 suggests that for ZP-LKE not only is the number of examples needed larger, but it also matters whether they are correct and known to the LLM (consistent with our hypothesis that examples are being used to infer the question). The two distinct ways in which examples are being used by FS-LKE and ZP-LKE map well to the dual modes of in-context learning namely, task recognition and task learning, respectively, that have been identified in a recent work(Pan, 2023).
We systematically investigate how factors such as how many examples are provided in an ZP-LKE, whether some of those examples are unknown to the model or simply incorrect, as well as how the examples are ordered affect knowledge estimation. We find that ZP-LKE requires many-shots, which make it relatively robust to unknown examples, but ZP-LKE remains vulnerable to incorrect examples. Our findings represent a nuanced exploration of in-context learning(Brown et al., 2020), where the dominant learning mode is task learning rather than task recognition.
3. ZP-LKE significantly outperforms previous prompt-based approaches across different open-source models and different types of factual relations: We empirically compared the performance of ZP-LKE against prior approaches that relied on a variety of human-generated prompts (HGPs) as well as machine-mined prompts (MMPs)(Jiang et al., 2020). Across a large set of facts spanning different types of relations from the widely-used T-Rex dataset (Elsahar et al., 2018), we find that ZP-LKE improves the fraction of facts accurately extracted from four open-source models by an average of 35% for HGPs (from 0.45 to 0.61) and 90% for MMPs (from 0.32 to 0.61). These performance gains of ZP-LKE arise from a better comprehension of the question as well as the expected answer format. To quantify the performance gains from only better question comprehension, we propose in Section 2.2 a multiple-choice testing that accounts for answer formats. We find that ZP-LKE still outperforms existing approaches by an average of 9.41% for HGPs (from 0.71 to 0.78) and 57% for MMPs (from 0.50 to 0.78), with improvements for specific relations like ”position played on team/specialty” varying from 152% for HGPs (from 0.17 to 0.43) and 310% for MMPs (from 0.10 to 0.43). Thus, ZP-LKE represents a better way to retrieve knowledge stored internally within an LLM, surpassing the model’s ability to follow instructions in prompt templates.
@@
4. Being model-agnostic, ZP-LKE enables a systematic comparison of latent knowledge of open source LLMs at scale: In contrast to prompt-based LKEs (Jiang et al., 2020; Shin et al., 2020), which are tailored to specific relations and models, ZP-LKE creates a single input to test for facts pertaining to a relation that can be used flexibly for any model. This simplicity and versatility allows for cross-LLM comparisions of factual knowledge. Using ZP-LKE, we evaluated the knowledge of 49 open-source LLMs from various families like Llama(2), Gemma, Mistral, OPT, and Pythia. These models vary in size and were tested with and without instruction-finetuning on 50 different relations and 20,000 facts from Wikidata. We found that models from families such as Llama2, Mistral, and Gemma, as well as larger models, know more facts. Models within the same family differ in the specific facts they know, even if trained on the same data. Additionally, instruction fine-tuning reduces the amount of factual knowledge that can be extracted from these models. Our findings will likely be of interest to developers that wish to train models with lots of embedded factual knowledge.
Related Work: Researchers have proposed several approaches (Youssef et al., 2023) to estimate latent knowledge from LLMs, which can be categorized into two main methods: (i) Model-internals based approaches: These methods use various internal aspects of the LLM, such as attention maps (Wang et al., 2020), activation functions (Burns et al., 2022), or model parameters (Kazemnejad et al., 2023), to determine whether factual information can be extracted from the model. (ii) Model-responses-based approaches, which are generally applicable to a wide range of LLM models and there are two key parts of the model-responses-based approach: constructing the input and evaluating the output from the LLM.
Input construction: There are different prompting techniques to verify if a target fact is stored in the model. These prompt-based methods differ in their choice of prompts, which can be divided into human-generated prompts (HGPs) (Cao et al., 2021; Chern et al., 2023; Sun et al., 2023; Wang et al., 2020; Petroni et al., 2019; Jiang et al., 2021; Newman et al., 2022; Jiang et al., 2020; Kalo and Fichtel, [n. d.]) and machine-mined prompts (MMPs) (Shin et al., 2020; Jiang et al., 2020). All the prompting-based methods try to find the best template for the question that the model can comprehend best; however, the optimization of the prompting template can be unreliable (Zamfirescu-Pereira et al., 2023; Arora et al., 2023; Sclar et al., 2023; Cao et al., 2021). Instead of finding the best comprehensive template, our approach proposes using structured triplet examples to communicate and prob the tested fact, uncover deeper relations and knowledge in the LLM. This method communicates the question through in-context examples, a strategy that, to our knowledge, has not been explored before. A similar approach is to use few-shot prompting. For example, (Kalo and Fichtel, [n. d.]) first found the best prompt template and then applied few-shot prompting to guide and limit the model’s response format. However, this approach still holds the limitation of template searching and relies on the model’s comprehension of the template, which is fundamentally different from our approach.
Output evaluation: The evaluation methods of early works are LLM specific, limiting the evaluated objects to single-token outputs (Bouraoui et al., 2020; Jiang et al., 2020; Shin et al., 2020; Petroni et al., 2019). More recent works evaluate the generation by checking the next generated tokens to see whether the potentially multi-token ground truth appears in the generated tokens (Cao et al., 2021; Chern et al., 2023; Sun et al., 2023; Wang et al., 2020; Jiang et al., 2021; Newman et al., 2022; Kalo and Fichtel, [n. d.]). However, the final performance is significantly influenced by the choice of , and the generation quality also relies heavily on various sampling parameters, which introduces uncertainty (Lin et al., 2023). In order to respond more fundamentally to the model’s level of knowledge, without focusing on metrics such as the fluency of generation, we constructed a multiple-choice dataset with 100 unique possible choices for each evaluated fact and judged whether or not the model knew the fact by comparing the relative probabilities of these 100 objects.
Factual knowledge datasets: Different from existing knowledge evaluation benchmarks like TruthfulQA (Lin et al., 2021) and MMLU (Hendrycks et al., 2020), which already provide templates of questions, our approach considers facts from existing knowledge graphs for performing knowledge estimation of LLMs. As a test bed (Elsahar et al., 2018; Hu et al., 2023a; Sun et al., 2023; Petroni et al., 2019; Zhu and Li, 2023; Kryściński et al., 2019), we utilize knowledge graphs, allowing our method to be applied to any knowledge graph database without additional effort.
2. Designing Reliable LKEs
@@
Today, there exist many general-purpose as well as domain-specific factual knowledge bases that contain a very large number (millions to billions) of facts. The facts can be encapsulated as triplets, represented as subject (), relation (), object (). These triplets offer a general way to represent factual knowledge about real-world entities in knowledge graphs or other structured knowledge bases. The goal of latent knowledge estimation is to infer what fraction of the facts are known to an LLM. We call methods that estimate the amount of latent knowledge inside an LLM latent knowledge estimators (LKEs).
2.1. Reliability concerns with existing LKEs
@@
Existing approaches to estimating latent knowledge in LLMs use various factual knowledge tests. We identify several reliability concerns (RCs) with current designs that motivate our new LKE design. While some related works address some of these concerns, none have comprehensively solved all the issues (Kalo and Fichtel, [n. d.]; Cao et al., 2021).
@@
RC 1. Reliance on unrestricted prompt engineering: Many past works have attempted to use test prompts without any restrictions, including both human-generated or machine-mined prompts (Jiang et al., 2020; Zamfirescu-Pereira et al., 2023; Arora et al., 2023; Sclar et al., 2023). They typically intersperse the subject and object between additional relationship context-communicating tokens. Some analyze the performance of a variety of prompts and then pick the best-performing or use an ensemble of the best-performing prompts (Jiang et al., 2020; Newman et al., 2022; Fernando et al., 2023). However, unrestricted prompt engineering risks introducing side-channels and over-fitting. First, the generated prompts, particularly those that are machine-mined, may include tokens that can implicitly or explicitly introduce additional (side-channel) information that makes it easier to answer the question. As a specific example, in a prior work (Jiang et al., 2020), for the relation “position held”, the prompt “ has the position of ” performed worse than “ is elected ”. But, note that the second prompt potentially introduces a side-channel: it implicitly rules out answer choices for unelected positions like Professor and favors elected positions like President. Second, selecting from an unbounded number of potential prompt choices raises concerns about the complexity of LKEs (the size of the set of all considered prompts) and the associated risks of over-fitting, which in turn affect the reliability of estimates.
@@
RC 2. Reliance on LLMs’ meta-linguistic judgments: Prior works used prompt templates with instructions (Chern et al., 2023; Sun et al., 2023; Wang et al., 2020; Petroni et al., 2019; Jiang et al., 2021; Newman et al., 2022; Jiang et al., 2020; Cao et al., 2021; Kalo and Fichtel, [n. d.]; Youssef et al., 2023; Zhao et al., 2024) for communicating the question as well as the expected format of answers. But, the scores (estimates) resulting from such prompt-based testing conflate an LLM’s latent knowledge of the facts with the LLM’s meta-linguistic judgments, i.e., the LLM’s ability to comprehend the prompt, understand the question embedded within the prompt and output the answer in some expected format (Hu and Levy, 2023). The impact on meta-linguistic judgments can be seen from the fact that multiple semantically-equivalent prompts result in different responses from an LLM and thereby, different estimates of latent knowledge (Hu and Levy, 2023).
RC 3. Reliance on LLM-specific prompts: Many prior works (Petroni et al., 2019; Jiang et al., 2020; Shin et al., 2020) limit the choice of facts that can be used in tests to those where the surface form of the objects () is represented by a single token by the LLM’s tokenizer. Even though some works are able to evaluate multiple-token objects, prompt-based approaches need careful prompt engineering for different LLMs to get the best prompt template (Cao et al., 2021; Chern et al., 2023; Sun et al., 2023; Wang et al., 2020; Jiang et al., 2021; Newman et al., 2022; Kalo and Fichtel, [n. d.]), which makes it hard to estimate and compare factual knowledge across a large number of LLMs and the prompt optimisation would be very expensive and inefficient for large models.
@@
Motivated by the above, we derive the following three design principles (DPs) for LKEs. A reliable LKE design should:
@@
DP1: limit prompt hacking to avoid over-fitting & side-channels.
DP2: minimize reliance on meta-linguistic prompts.
DP3: avoid LLM-specific prompts.
2.2. A new Zero- Prompt based LKE (ZP-LKE)